OpenStudy (anonymous):

If the area of a rectangle 16s2t and the length is 8st2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer.

OpenStudy (anonymous):

2st 2/st 2s/t t/2s

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

nevermind, it's C

OpenStudy (jdoe0001):

\(\bf Area = Length\cdot Width\qquad Area=16s^2t\qquad Length=8st^2 \\ \quad \\ Area = Length\cdot Width\implies \cfrac{Area}{Length}=Width\implies \cfrac{16s^2t}{8st^2d}=Width\)